OpenStudy (anonymous):

Three men went to a hotel to rent a room, the cost of the room was $30. Each man paid $10 to the bellboy and proceeded to their room. After a little while the bellboy realized that there was a special on rooms that night and the price for the men's room should have been $25. On his way to the men's room to give them back $5, he was puzzled how he was going to split $5 as he had no change. He decided he would give them each $1 and keep the remaining $2 for himself. So each man originally paid $10, but after the bellboy gave each man $1 back, each man paid $9. 9 x 3 = $27 plus the $2 the bellbo

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

the other half is missing

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

it's already too long without the missing half :)

OpenStudy (anonymous):

put in his pocket equals $29. The original price for the room was $30. Where did the last dollar go?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

They paid $30, then the bellboy took $5 to give back. 30-5=25. The bellboy gave each of them $1; 1*3 = 3 25+3=28 The bellboy kept $2; 28+2 = 30.

OpenStudy (amistre64):

yep; since the room was $25 they overpaid by $3..gotta track down the bboy for the extra $2